# Valid Conformal Prediction for Dynamic GNNs

### v1

@@ Abstract
Graph neural networks (GNNs) are powerful black-box models which have shown impressive empirical performance. However, without any form of uncertainty quantification, it can be difficult to trust such models in high-risk scenarios. Conformal prediction aims to address this problem, however, an assumption of exchangeability is required for its validity which has limited its applicability to static graphs and transductive regimes.
We propose to use unfolding, which allows any existing static GNN to output a dynamic graph embedding with exchangeability properties. Using this, we extend the validity of conformal prediction to dynamic GNNs in both transductive and semi-inductive regimes. We provide a theoretical guarantee of valid conformal prediction in these cases and demonstrate the empirical validity, as well as the performance gains, of unfolded GNNs against standard GNN architectures on both simulated and real datasets.
@@ 1 Introduction
Graph neural networks (GNNs) have seen formidable success in a wide variety of application domains including prediction of protein structure [1] and molecular properties [2, 3], drug discovery [4], computer vision [5], natural language processing [6, 7], recommendation systems [8, 9, 10], estimating time of arrival (ETA) in services like Google Maps [11], and advancing mathematics [12]. They often top leaderboards in benchmarks relating to machine-learning on graphs [13], for a range of tasks such as node, edge, or graph property prediction. Useful resources for getting started with GNNs include the introductions [14, 15] and the Pytorch Geometric library [16].
Conformal prediction, advanced by Vovk and co-authors [17], and later expanded on by various researchers in the Statistics community [18, 19, 20, 21, 22, 23, 24, 25, 26], is an increasingly popular paradigm for constructing provably valid prediction sets from a ‘black-box’ algorithm with minimal assumptions and at low cost. Several recent papers have brought this idea to quantifying uncertainty in GNNs [27, 28, 29].
There are many applications of significant societal importance, e.g. cyber-security [30, 31, 32], human-trafficking prevention [33, 34], social media & misinformation [35], contact-tracing [36], patient care [37, 38] where we would like to be able to use GNNs, with uncertainty quantification, on dynamic graphs.
The added time dimension in dynamic graphs causes serious issues with existing GNN + conformal methodology, broadly relating to de-alignment between embeddings across time points (see Figure 1), resulting in large or invalid prediction sets in even simple and stable dynamic regimes.
Inspired by a concurrent line of statistical research on spectral embedding and tensor decomposition [39, 40, 41, 42, 43, 44, 45], we propose to use one of the unfoldings [46] of the tensor representation of the dynamic graph as input to a standard GNN, and demonstrate validity in both transductive and semi-inductive regimes.
@@
The only paper we found studying conformal inference on dynamic graphs is [29], but the paper considers a strictly growing graph, which is not what we mean by a ‘dynamic graph’, where edges appear and disappear in complex and random ways, as in all of our examples and applications alluded to above. Our work naturally builds on earlier contributions combining GNNs and conformal inference [27, 28, 29] but our theory is distinct from those contributions, with arguments closer to [24], and the use of unfoldings for GNNs is new. In the dynamic graph literature, unfoldings have exclusively been proposed for unsupervised settings [43, 44, 45, 47, 48], particularly spectral embedding, and have never been applied to conformal inference.
@@ 2 Theory & Methods
Problem setup. In this paper, a dynamic graph $G$ is a sequence of $T$ graphs over a globally defined nodeset $[n]=\{1,\ldots,n\}$, represented by adjacency matrices $\mathbf{A}^{(1)},\ldots,\mathbf{A}^{(T)}$. It may help to assume the $\mathbf{A}^{(t)}$ are $n\times n$ binary symmetric matrices corresponding to undirected graphs, but in fact all we require is that the $\mathbf{A}^{(t)}$ have the same number of rows, $n_{r}$, allowing for directed, bipartite and weighted graphs.
@@
We track the labelled and unlabelled pairs in this problem using a fixed sequence $\nu$ where for each $(i,t)\in\nu$ we assume existence of
@@
    a class label $Y^{(t)}_{i}\in\mathcal{Y}$, which may or may not be observed, where $\mathcal{Y}$ is some discrete set of cardinality $d$;
@@
    a set of attributes, which is observed, and represented by a feature vector $X^{(t)}_{i}\in\mathbb{R}^{p}$;
@@
    a corresponding column in $\mathbf{A}^{(t)}$, denoted $\mathbf{A}^{(t)}_{i}$.
@@
Transductive regime. For some $m<|\nu|$, the missing label is missing completely at random among the first $m+1$ node/time pairs of $\nu$.
Semi-inductive regime. For some $m<|\nu|$, the missing label corresponds to a fixed pair among the first $m+1$ node/time pairs of $\nu$.
Let $\nu_{\text{test}}$ denote the missing node/time pair, $\nu^{(\text{training})}\coloneqq\{\nu_{\ell},\ell>m+1\}$, and $\nu^{(\text{calibration})}\coloneqq\{\nu_{\ell},\ell\leq m+1\}\setminus\{\nu_{\text{test}}\}$.
In the transductive regime, a valid confidence set can be constructed with no further assumptions (Algorithm 1). This may seem surprising and could be very useful in view of the following implication: the pairs $\nu^{(\text{training})}$ can correspond to historical datapoints, at times $t<T$, and the pairs $\{\nu_{\text{test}}\}\cup\nu^{(\text{calibration})}$ can correspond to datapoints in the present, $T$. We only need the missing label to be uniformly chosen among the labels at time $T$. The semi-inductive regime is more challenging, and could reflect a scenario where we have only historical labels. Here, we will need to invoke further assumptions such as symmetry and exchangeability, familiar in the conformal prediction literature.
Proposed approach. We will repurpose a standard GNN, designed for static graphs, into producing dynamic graph embeddings with desirable exchangeability properties. Let $\mathcal{G}$ denote a GNN taking as input an adjacency matrix $\mathbf{A}$ corresponding to an undirected graph on $N$ nodes, along with node attributes and labels (some missing), and returning an embedding $\hat{\mathbf{Y}}\in\mathbb{R}^{N\times d}$. Typically, the $i$th row of $\hat{\mathbf{Y}}$ predicts the label of node $i$ through
^(node i has label $\hat{\mathbb{P}}(\text{node $i$ has label $k$})=\left[\text{softmax}(\hat{\mathbf{Y}}_{i1},\ldots,\hat{\mathbf{Y}}_{id})\right]_{k}.$
In place of this generic input $\mathbf{A}$, we will enter the dilated unfolding [47] corresponding to the dynamic graph $G$,
$\mathbf{A}^{\text{UNF}}\coloneqq\begin{bmatrix}\mathbf{0}&\mathbfcal{A}\\ \mathbfcal{A}^{\top}&\mathbf{0}\end{bmatrix},$
@@
The resulting output from the GNN splits into embeddings
@@
where $\hat{\mathbf{X}}^{\text{UNF}}\in\mathbb{R}^{n_{r}\times d}$ contains global representations. The embedding $\hat{\mathbf{Y}}^{\text{UNF}}$ contains representations of node/time pairs, of principal interest here, and we let $\hat{\mathbf{Y}}^{\text{UNF}}_{w}$ denote the row representing node/pair $w\in\nu$.
Finally, we shall require some non-conformity score function $r:\mathbb{R}^{d}\times\mathcal{Y}\rightarrow\mathbb{R}$, with the convention that $r(\hat{y},y)$ is large when $\hat{y}$ is a poor prediction of $y$. In practice, we make use of the adaptive non-conformity measures detailed by [26].
With these ingredients in place, we are in a position to obtain a prediction set $\hat{C}$ for $\nu_{\text{test}}$, the computation of which is detailed in Algorithm 1. The notation leaves implicit the dependence of the set $\hat{C}$ on the observed data; explicitly, the set $\hat{C}$ is a deterministic function of the dynamic graph $G$, the attributes $(X_{w},w\in\nu)$, and the labels $(Y_{w},w\in\nu\setminus\{\nu_{\text{test}}\})$.
@@
Input: Dynamic graph $G$, node/time pairs $\nu=\nu^{(\text{training})}\cup\nu^{(\text{calibration})}\cup\{\nu_{\text{test}}\}$, attributes $(X_{w},w\in\nu)$, labels $(Y_{w},w\in\nu\setminus\{\nu_{\text{test}}\})$, confidence level $\alpha$, non-conformity function $r$
1:Learn $\hat{\mathbf{Y}}^{\text{UNF}}$ based on $\nu^{(\text{training})}$
2:Let $\hat{C}=\mathcal{Y}$
3:Let $R_{w}=r(\hat{\mathbf{Y}}^{\text{UNF}}_{w},Y_{w})$ for $w\in\nu^{(\text{calibration})}$
@@
$\hat{q}=\text{$\lfloor\alpha(m+1)\rfloor$ largest value in $R_{w},w\in\nu^{(\text{calibration})}$}$
@@
6:     Let $R_{\text{test}}=r(\hat{\mathbf{Y}}^{\text{UNF}}_{\nu_{\text{test}}},y)$
@@
$\mathbb{P}(Y_{\nu_{\text{\rm test}}}\in\hat{C})\geq 1-\alpha.$
Let $\nu^{+}_{\text{calibration}}=(\nu_{\ell},\ell\in[m+1])$, comprising $\nu_{\text{calibration}}$ and $\nu_{\text{test}}$. For the semi-inductive regime, we make the following assumptions:
@@
The columns $\mathbf{A}_{w}$, attributes $X_{w}$, and labels $Y_{w}$ corresponding to $w\in\nu^{+}_{\text{\rm calibration}}$ are jointly exchangeable.
We make the definition of exchangeability precise in the supplementary material. Informally, we can swap columns of $\mathcal{A}=\left(\mathbf{A}^{(1)},\dots,\mathbf{A}^{(T)}\right)$ corresponding to test or calibration pairs, along with the attributes and labels, without changing the likelihood of the data.
@@
The GNN $\mathcal{G}$ is label equivariant.
This definition is again made precise in the supplementary material. Informally, if we re-order the nodes, attributes and labels, we re-order the resulting embedding.
@@
$\mathbb{P}(Y_{\nu_{\text{\rm test}}}\in\hat{C})\geq 1-\alpha.$
Discussion. Assumption 2 is relatively mild and roughly satisfied by standard GNN architectures. Assumption 1 is more challenging, and can guide us towards curating a calibration set. For example, if we hypothesise that the $(\mathbf{A}^{(t)}_{i},t\in[T])$ are i.i.d. stationary stochastic processes (over time), then we can satisfy Assumption 1 by ensuring $\nu^{+}_{\text{calibration}}$ contains only distinct nodes. The independence hypothesis will typically break in undirected graphs because of the necessary symmetry of each $\mathbf{A}^{(t)}$, and in this case we may also wish to ensure that $\nu^{+}_{\text{calibration}}$ contains only distinct time points. In general, for predicting the label of nodes based on historical information, Assumption 1 should make us wary of global drift, a poor distributional overlap between the test and calibration nodes, and autocorrelation. The semi-inductive regime is where our theory actively requires a dilated unfolding, and where alternative approaches can fail completely; however, even in the transductive regime where our theory covers alternative approaches, the prediction sets from unfolded GNNs (UGNNs) are often smaller.
@@
Consider a simple two-community dynamic stochastic block model (DSBM) [49, 50] for undirected graphs. Let
@@
A typical way of applying a GNN to a series of graphs is to first stack the adjacency matrices as blocks on the diagonal [51],
@@
$\hat{\mathbf{Y}}_{\text{BD}}=\mathcal{G}\left(\mathbf{A}^{\text{BD}}\right).$
Figure 1 plots the representations, $\hat{\mathbf{Y}}_{\text{BD}}$ and $\hat{\mathbf{Y}}_{\text{UNF}}$ of $\mathbf{A}^{(1)},\mathbf{A}^{(2)}$ using a GCN [52] as $\mathcal{G}$. It is clear that, despite the fact that $\mathbf{A}^{(1)}$ and $\mathbf{A}^{(2)}$ come from the same distribution, the output of the block diagonal GCN appears to encode a change that is not present. We refer to this issue as temporal shift. In contrast, the output of UGCN is exchangeable over time and so features no temporal shift.
Figure 1: UGCN and block diagonal GCN representations of an i.i.d. stochastic block model after applying PCA. The models were trained with transductive masks. Block diagonal GCN appears to encode a significant change over time despite there being none. The embedding from UGCN is exchangeable over time, as would be expected.

### v2

@@ Abstract
Dynamic graphs provide a flexible data abstraction for modelling many sorts of real-world systems, such as transport, trade, and social networks. Graph neural networks (GNNs) are powerful tools allowing for different kinds of prediction and inference on these systems, but getting a handle on uncertainty, especially in dynamic settings, is a challenging problem.
In this work we propose to use a dynamic graph representation known in the tensor literature as the unfolding, to achieve valid prediction sets via conformal prediction. This representation, a simple graph, can be input to any standard GNN and does not require any modification to existing GNN architectures or conformal prediction routines.
One of our key contributions is a careful mathematical consideration of the different inference scenarios which can arise in a dynamic graph modelling context. For a range of practically relevant cases, we obtain valid prediction sets with almost no assumptions, even dispensing with exchangeability. In a more challenging scenario, which we call the semi-inductive regime, we achieve valid prediction under stronger assumptions, akin to stationarity.
We provide real data examples demonstrating validity, showing improved accuracy over baselines, and sign-posting different failure modes which can occur when those assumptions are violated.
@@ 1 Introduction
Graph neural networks (GNNs) have seen success in a wide variety of application domains including prediction of protein structure [Jumper et al., 2021] and molecular properties [Duvenaud et al., 2015, Gilmer et al., 2017], drug discovery [Stokes et al., 2020], computer vision [Sarlin et al., 2020], natural language processing [Peng et al., 2018, Wu et al., 2023], recommendation systems [Berg et al., 2017, Wu et al., 2022, Borisyuk et al., 2024], estimating time of arrival (ETA) in services like Google Maps [Derrow-Pinion et al., 2021], and advancing mathematics [Davies et al., 2021]. They often top leaderboards in benchmarks relating to machine-learning on graphs [Ope, ], for a range of tasks such as node, edge, or graph property prediction. Useful resources for getting started with GNNs include the introductions [Hamilton, 2020, Sanchez-Lengeling et al., 2021] and the Pytorch Geometric library [Pyt, ].
Conformal prediction (CP), advanced by Vovk and co-authors [Vovk et al., 2005], and later expanded on by various researchers in the Statistics community [Shafer and Vovk, 2008, Lei et al., 2013, Lei and Wasserman, 2014, Lei et al., 2018, Foygel Barber et al., 2021, Tibshirani et al., 2019, Barber et al., 2023, Gibbs et al., 2023, Romano et al., 2020], is an increasingly popular paradigm for constructing provably valid prediction sets from a ‘black-box’ algorithm with minimal assumptions and at low cost. Several recent papers have brought this idea to quantifying uncertainty in GNNs [Huang et al., 2024b, Clarkson, 2023, Zargarbashi and Bojchevski, 2023].
There are many applications of significant societal importance, e.g. cyber-security [Bilot et al., 2023, He et al., 2022, Bowman and Huang, 2021], human-trafficking prevention [Szekely et al., 2015, Nair et al., 2024], social media & misinformation [Phan et al., 2023], contact-tracing [Tan et al., 2022], patient care [Liu et al., 2020, Boll et al., 2024] where we would like to be able to use GNNs, with uncertainty quantification, on dynamic graphs.
The added time dimension in dynamic graphs causes serious issues with existing GNN + conformal methodology, broadly relating to de-alignment between embeddings across time points (see Figure 1), resulting in large or invalid prediction sets in even simple and stable dynamic regimes. The unsolved problem we address, in a nutshell, is guaranteeing that two nodes behaving in the same way at distinct points in time are embedded in an exchangeable way, regardless of global network dynamics.
Inspired by a concurrent line of statistical research on spectral embedding and tensor decomposition [Zhang and Xia, 2018, Cape et al., 2019, Abbe et al., 2020, Rubin-Delanchy et al., 2022, Jones and Rubin-Delanchy, 2020, Gallagher et al., 2021, Agterberg and Zhang, 2022], we propose to use one of the unfoldings [De Lathauwer et al., 2000] of the tensor representation of the dynamic graph as input to a standard GNN, and demonstrate validity in both transductive and semi-inductive regimes.
@@
The use of (tensor) unfoldings in GNNs is new. At present, to our knowledge, the only paper studying conformal inference on dynamic graphs is [Zargarbashi and Bojchevski, 2023], but the paper considers a strictly growing graph, which is very different from our meaning, in which a ‘dynamic graph’ has edges which appear and disappear in complex and random ways, as in all of our examples and in the applications alluded to above. The motivation for our work derives from earlier contributions studying CP on static graphs using GNNs [Huang et al., 2024b, Clarkson, 2023, Zargarbashi and Bojchevski, 2023] but our theory is distinct from those contributions, with arguments closer to [Barber et al., 2023]. In the dynamic graph literature, unfoldings have exclusively been proposed for unsupervised settings [Jones and Rubin-Delanchy, 2020, Gallagher et al., 2021, Agterberg and Zhang, 2022, Davis et al., 2023, Wang et al., 2023], particularly spectral embedding, and have never been applied to conformal inference.
Our contribution should be viewed as a novel interface between CP and GNNs, rather than an improvement of either. This interface allows for a modular system design in dynamic graph inference in which both the GNN and CP modules can be chosen and optimised independently. In view of the rapid development of both fields, we have opted against fine-tuning to allow more transparent, stable, and reproducible performance assessments using recognised and standard baselines for GNNs and CP: Graph Convolutional Networks (GCN) [Kipf and Welling, 2016] and Graph Attention Networks (GAT) [Veličković et al., 2017] for GNNs and APS [Romano et al., 2020] for CP. Our theory applies to the representation of the dynamic graph, not the choice of GNN or CP procedures.
Figure 1: Contribution overview. a) This paper is about the representation of the collection of adjacency matrix snapshots. The baseline (current practice) approach treats these as independent and can be viewed as padding a ‘block-diagonal’ matrix with zeroes. Unfolding instead column concatenates which links nodes to themselves over time. Dilation results in a square symmetric matrix. b) Which data are available at training time affects performance; we report results for transductive (all time-points are exchangeable in terms of test/train split), semi-inductive (a future period is reserved for testing), or temporal transductive (a future period is reserved for testing and calibration). c) Simulation of an i.i.d. stochastic block model showing the embedding after applying PCA. The models were trained with transductive masks. Block diagonal GCN appears to encode a significant change over time despite there being none. The embedding from UGCN is exchangeable over time, as would be expected.
@@ 2 Theory & Methods
Problem setup. In this paper, a dynamic graph $G$ is a sequence of $T$ graphs over a globally defined nodeset $[n]=\{1,\ldots,n\}$, represented by adjacency matrices $\mathbf{A}^{(1)},\ldots,\mathbf{A}^{(T)}$. Technically, all we require is that the $\mathbf{A}^{(t)}$ have the same number of rows, $p$, allowing for directed, bipartite and weighted graphs. However, the reader may find it easier to assume the $\mathbf{A}^{(t)}$ are $n\times n$ binary symmetric matrices corresponding to undirected graphs, in which case $p=n$.
@@
We track the labelled and unlabelled node/time pairs in this problem using a fixed sequence $\nu$ where for each $(i,t)\in\nu$ we assume existence of
@@
    a class label $Y^{(t)}_{i}\in\mathcal{Y}$, which may not be observed, in some discrete set of cardinality $d$;
@@
    a set of attributes, which is observed, and represented by a feature vector $X^{(t)}_{i}\in\mathbb{R}^{c}$;
@@
    a corresponding column in $\mathbf{A}^{(t)}$, denoted $\mathbf{A}^{(t)}_{i}$, with at least one non-zero entry.
@@
The data are split into four sets: $\nu^{(\text{training})}$ to learn the GNN parameters; $\nu^{(\text{validation})}$ for out-of-sample performance evaluation of the GNN; $\nu^{(\text{calibration})}$ to train the conformal model, and $\nu^{(\text{test})}$ is the target for conformal inference.
Clearly, the theory and methodology of this paper are applicable to multiple test nodes, but for simplicity in this section we assume $\nu^{(\text{test})}$ comprises only one node.
From these objects we define three versions of the inference problem.
Transductive regime. The test label is missing completely at random allowing training, validation, and calibration at any time (Figure 1b, top).
Temporal transductive regime. Training and validation are undertaken up to some time $\tau$. The test label is missing completely at random after time $\tau$, allowing calibration from this interval (Figure 1b, bottom).
Semi-inductive regime. Training and validation is undertaken up to some time $\tau$. The test label occurs after time $\tau$, corresponding to a fixed node $i$, preventing calibration from this interval (Figure 1b, middle).
In either of the transductive regimes, a valid confidence set can be constructed with no further assumptions (Algorithm 1). We believe this could be exciting because we view the temporal transductive regime, in particular, as an extremely common scenario. The semi-inductive regime is more challenging, and reflects a scenario where we have only historical labels. Here, we will need to invoke further assumptions such as symmetry and exchangeability, familiar in the conformal prediction literature.
Formally, for some $m<|\nu|-1$, we reorder the dataset $\nu$ so that the first $m$ node/time pairs correspond to the calibration and test sets, while the remaining pairs correspond to the training and validation sets. For both transductive regimes, the missing label is missing at random among the first $m$ node/time pairs of $\nu$; for the semi-inductive regime, the missing label corresponds to a fixed node/time pair in this set. Let $\nu^{(\text{test})}\coloneqq\{w_{\text{test}}\}$ denote the missing node/time pair, $\nu^{(\text{calibration})}\coloneqq\{w_{\ell},\ell\leq m\}\setminus\nu^{(\text{test})}$, and $\nu^{(\text{training})}\cup\nu^{(\text{validation})}\coloneqq\{w_{\ell},\ell>m\}$.
Proposed approach. We will repurpose a standard GNN, designed for static graphs, to produce dynamic graph embeddings with desirable exchangeability properties. Let $\mathcal{G}$ denote a GNN taking as input an adjacency matrix corresponding to an undirected graph on $n$ nodes, along with node attributes and node labels, and returning an embedding $\hat{\mathbf{V}}\in\mathbb{R}^{n\times d}$. Typically, the $i$th row of $\hat{\mathbf{V}}$ predicts the label of node $i$ through
^(node i has label $\hat{\mathbb{P}}(\text{node $i$ has label $k$})=\left[\text{softmax}(\hat{\mathbf{V}}_{i1},\ldots,\hat{\mathbf{V}}_{id})\right]_{k}.$
To represent the adjacency matrix we apply dilated unfolding [Davis et al., 2023] representing the dynamic graph $G$ in the form
$\mathbf{A}^{\text{UNF}}\coloneqq\begin{bmatrix}\mathbf{0}&\mathbfcal{A}\\ \mathbfcal{A}^{\top}&\mathbf{0}\end{bmatrix}\in\mathbb{R}^{(p+nT)\times(p+nT)}$
@@
The GNN takes as input a matrix $X^{\text{UNF}}\in\mathbb{R}^{(p+nT)\times c}$ containing $c$-dimensional attributes for all nodes in the dilated unfolding. For networks without attributes this matrix is the $(p+nT)\times(p+nT)$ identity matrix. The GNN also requires training and validation labels $Y_{w},w\in\nu^{(\text{training})}\cup\nu^{(\text{validation})}$, where global nodes are not assigned a label. The resulting output splits into embeddings
@@
where $\hat{\mathbf{U}}^{\text{UNF}}\in\mathbb{R}^{p\times d}$ contains global representations and the embedding $\hat{\mathbf{V}}^{\text{UNF}}\in\mathbb{R}^{nT\times d}$ contains a representation of each node/time pair, of principal interest here.
Finally, we shall require some non-conformity score function $r:\mathbb{R}^{d}\times\mathcal{Y}\rightarrow\mathbb{R}$, with the convention that $r(\hat{y},y)$ is large when $\hat{y}$ is a poor prediction of $y$. In practice, we make use of the adaptive non-conformity measures detailed by Romano et al. [2020].
With these ingredients in place, we are in a position to obtain a prediction set $\hat{C}$ for $w_{\text{test}}$, the computation of which is detailed in Algorithm 1. The notation leaves implicit the dependence of the set $\hat{C}$ on the observed data; explicitly, the set $\hat{C}$ is a deterministic function of the dynamic graph $G$, the attributes $(X_{w},w\in\nu)$, and the labels $(Y_{w},w\in\nu\setminus\nu^{(\text{test})})$.
@@
Input: Dynamic graph $G$, node/time pairs $\nu=\nu^{(\text{training})}\cup\nu^{(\text{validation})}\cup\nu^{(\text{calibration})}\cup\nu^{(\text{test})}$, attributes $(X_{w},w\in\nu)$, labels $(Y_{w},w\in\nu\setminus\nu^{(\text{test})})$, confidence level $\alpha$, non-conformity function $r$
1:Learn $\hat{\mathbf{V}}^{\text{UNF}}$ based on $\nu^{(\text{training})}$ and $\nu^{(\text{validation})}$
2:Let the initial confidence set be $\hat{C}=\mathcal{Y}$
3:Let $R_{w}=r(\hat{\mathbf{V}}^{\text{UNF}}_{w},Y_{w})$ for $w\in\nu^{(\text{calibration})}$
@@
$\hat{q}=\text{$\lfloor\alpha(m)\rfloor$ largest value in $R_{w},w\in\nu^{(\text{calibration})}$}$
@@
6:     Let $R_{\text{test}}=r(\hat{\mathbf{V}}^{\text{UNF}}_{w_{\text{test}}},y)$
@@
$\mathbb{P}(Y_{w_{\text{\rm test}}}\in\hat{C})\geq 1-\alpha.$
For the semi-inductive regime, we make the following assumptions:
@@
The columns $\mathbf{A}_{w},w\in\nu$, attributes $X_{w},w\in\nu$, and labels $Y_{w},w\in\nu$ are jointly exchangeable, that is, for any permutation $\pi:[m]\rightarrow[m]$,
The symbol $\triangleq$ means “equal in distribution”. Informally, we can swap columns of $\mathcal{A}=\left(\mathbf{A}^{(1)},\dots,\mathbf{A}^{(T)}\right)$ corresponding to test or calibration pairs, along with the attributes and labels, without changing the likelihood of the data.
Given a permutation $\pi:[m]\rightarrow[m]$, we use $\pi G$ to denote a dynamic graph in which the columns of $G$ corresponding to $\nu_{1},\ldots,\nu_{m}$ are permuted according to $\pi$. More precisely, $\pi G$ is the dynamic graph corresponding to the sequence of adjacency matrices $\mathbf{A}^{(1)^{\prime}},\ldots,\mathbf{A}^{(T)^{\prime}}$ obtained as follows. First, copy $\mathbf{A}^{(1)},\ldots,\mathbf{A}^{(T)}$ to $\mathbf{A}^{(1)^{\prime}},\ldots,\mathbf{A}^{(T)^{\prime}}$. Next, overwrite any column $\nu_{\ell}\in\nu_{1},\ldots,\nu_{m}$ to $\mathbf{A}^{\prime}_{\nu_{\ell}}=\mathbf{A}_{\nu_{\pi^{-1}(\ell)}}$. If $x$ is a sequence of length at least $m$, we write $\pi x$ to denote the same sequence with its first $m$ elements permuted according to $\pi$, that is, $(\pi x)_{\ell}=x_{\pi^{-1}(\ell)}$ if $\ell\leq m$ and $(\pi x)_{\ell}=x_{\ell}$ otherwise.
@@
The GNN $\mathcal{G}$ is permutation equivariant, that is, for any permutation $\pi:[N]\rightarrow[N]$ with associated permutation matrix $\Pi$,
$\mathcal{G}(\Pi\mathbf{A}\Pi^{\top})=\Pi\mathcal{G}(\mathbf{A}).$
Informally, if we re-order the nodes, attributes and labels, we re-order the resulting embedding.
@@
$\mathbb{P}(Y_{w_{\text{\rm test}}}\in\hat{C})\geq 1-\alpha.$
Discussion. Assumption 2 is relatively mild and roughly satisfied by standard GNN architectures. Assumption 1 is more challenging, and can guide us towards curating a calibration set. For example, if we hypothesise that the $(\mathbf{A}^{(t)}_{i},t\in[T])$ are i.i.d. stationary stochastic processes (over time), then we can satisfy Assumption 1 by ensuring $\nu^{(\text{calibration})}\cup\nu^{(\text{test})}$ contains only distinct nodes. The independence hypothesis will typically break in undirected graphs because of the necessary symmetry of each $\mathbf{A}^{(t)}$, and in this case we may also wish to ensure that $\nu^{(\text{calibration})}\cup\nu^{(\text{test})}$ contains only distinct time points. In general, for predicting the label of nodes based on historical information, Assumption 1 should make us wary of global drift, a poor distributional overlap between the test and calibration nodes, and autocorrelation. The semi-inductive regime is where our theory actively requires a dilated unfolding, and where alternative approaches can fail completely; however, even in the transductive regime where our theory covers alternative approaches, the prediction sets from unfolded GNNs (UGNNs) are often smaller.
@@
Consider a simple two-community dynamic stochastic block model (DSBM) [Yang et al., 2011, Xu and Hero, 2014] for undirected graphs. Let
@@
A typical way of applying a GNN to a series of graphs is to first stack the adjacency matrices as blocks on the diagonal [Fey and Lenssen, 2019],
@@
$\hat{\mathbf{V}}^{\text{BD}}=\mathcal{G}\left(\mathbf{A}^{\text{BD}},X^{\text{BD}},\{Y_{w}:w\in\nu^{(\text{training})}\cup\nu^{(\text{validation})}\}\right),$
where $X^{\text{BD}}\in\mathbb{R}^{nT\times c}$ is the matrix of $c$-dimensional attributes for all node/time pairs. (For networks without attributes, $X^{\text{BD}}$ is set to be the $nT\times nT$ identity matrix). From this representation, it is clear that $\mathbf{A}^{\text{BD}}$ is not exchangeable over time.
Figure 1 plots the representations, $\hat{\mathbf{V}}_{\text{BD}}$ and $\hat{\mathbf{V}}_{\text{UNF}}$ of $\mathbf{A}^{(1)},\mathbf{A}^{(2)}$ using a GCN [Kipf and Welling, 2016] as $\mathcal{G}$. It is clear that, despite the fact that $\mathbf{A}^{(1)}$ and $\mathbf{A}^{(2)}$ come from the same distribution, the output of the block diagonal GCN appears to encode a change that is not present. We refer to this issue as temporal shift. In contrast, the output of UGCN is exchangeable over time and so features no temporal shift.